OpenStudy (anonymous):

simplify 1) -17^0 +(-17)^0 what are the rules to adding exponents and dividing them

OpenStudy (anonymous):

anything to the power of zero is one, so this simplifies to -1+ -1 = -1-1 = -2

OpenStudy (anonymous):

I believe it is 1+1=2.

OpenStudy (lgbasallote):

it's actually -1 + 1 ^_^ -(17^0) + (-17)^0 = -(1) + (1) = -1 + 1

OpenStudy (anonymous):

I made a mistake the answer should be 0, @lgbasallote is right!

OpenStudy (anonymous):

ok thnks but what sre the rules i asked for

OpenStudy (lgbasallote):

oh..uhmmm you add exponents when you have the same base. for example: \(\large a^m \times a^n = a^{(m+n)}\) you divide exponents when you have a radical and an exponent..for example.. \(\large \sqrt[3]{a^2} = a^{\frac{2}{3}}\) make sense?

OpenStudy (lgbasallote):

adding exponents can only be done when multiplying two numbers/letters that have the same base

OpenStudy (anonymous):

yeah but what if you have hold on ill draw it for you

OpenStudy (anonymous):